OpenStudy (anonymous):

Will medal+fan PLEASE HELP!! Can someone help and explain how to solve this question. Write y=(-5/8)x+3 in standard form using integers.

OpenStudy (anonymous):

@uri @nincompoop @Mehek14 @mathmale @pooja195

jhonyy9 (jhonyy9):

do you know what mean ,,standard form" of y ?

jhonyy9 (jhonyy9):

you need to know formula of ...

jhonyy9 (jhonyy9):

y=mx+c - yes ?

OpenStudy (anonymous):

ya

jhonyy9 (jhonyy9):

how you can making integers from this fraction ?

OpenStudy (anonymous):

I don't know

OpenStudy (mrnood):

(y=mx+c) is NOT the standard form that is known as the slope intercept form 'Standard' form is Ax + By = C start by multiplying both sides of equation by 8

OpenStudy (anonymous):

ok but y 8?

OpenStudy (mrnood):

to get rid of the fraction (-5/8) that will produce an equation with just integers in it. You then need to rearrange it into the form I wrote above: Ax + By = C

OpenStudy (anonymous):

oh okay so 8* y=(-5/8)x+3 *8= 8*y=(-5)x+3 right?

OpenStudy (mrnood):

no - you didn't multiply properly 8y = 8(-5x/8 +3) multiply ALL the terms in the brackets

OpenStudy (anonymous):

3 isn't in the brackets tho

OpenStudy (mrnood):

the right side of the equation is -5/8 x + 3 you need to multiply the whole of that by 8 so I put the brackets there to make it clearer

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

so imultiply -5 and 3 by 8?

OpenStudy (mrnood):

you started with y = -5/8 x + 3 to get rid of the fraction you need to multiply the WHOLE equation by 8 so 8(y) = 8(-5/8 x +3) so multiply the 8 into the brackets

OpenStudy (anonymous):

so -5x+24

OpenStudy (mrnood):

8y = -5x + 24 is the correct expansion but to get to standard form you now need to ADD 5x to each side...

OpenStudy (anonymous):

-5x or just 5x

OpenStudy (mrnood):

Add 5x that will get the constant term on its own on the right side...

OpenStudy (anonymous):

ok so 8y+5x=-5x+3+5x right?

OpenStudy (mrnood):

no - you have copied wrong... the left side is correct (and then simplify the right side)

OpenStudy (anonymous):

how?

OpenStudy (mrnood):

first : you wrote 3 where there should be 24 8y +5x = 5x -5x +24 simplify that

OpenStudy (anonymous):

8y+5x=5x+24?

OpenStudy (mrnood):

8y +5x = (5x -5x) +24 simplify that

OpenStudy (anonymous):

8y+5x=x+24

OpenStudy (mrnood):

no come on - look at the equation we have done all this and you simply need to simplify it so it looks like Ax+By = C

OpenStudy (mrnood):

8y +5x = (5x -5x) +24 (hint - what is (5x-5x))

OpenStudy (anonymous):

ok so iknow i've been so stupid and uve so paitent (thank you :) ) but i have another guess 8y+5x=1 i think

OpenStudy (mrnood):

don't GUESS that is not correct hint what is 1-1 what is 2-2 what is 3-3 what is 5x-5x

OpenStudy (anonymous):

ohhh its 0 x) i'm so stupid so 8y+5x=0

OpenStudy (mrnood):

what happened to the +24?

OpenStudy (anonymous):

8y+5x=24

OpenStudy (mrnood):

That's correct - well done so what we did was: first get rid of the fraction by multiplying by the denominator THEN arrange the equation by moving the x and y terms to one side, leaving the constant term(24) on the other

OpenStudy (anonymous):

FINALLY thanks sooooo much :D

OpenStudy (mrnood):

np